### v1

@@ Abstract
In this paper, we report the detection of the very-high-energy (VHE, $100{\rm\ GeV}<E<100{\rm\ TeV}$) and ultra-high-energy (UHE, $E>100\rm\ TeV$) $\gamma$-ray emissions from the direction of the young star-forming region W43, observed by the Large High Altitude Air Shower Observation (LHAASO). The extended $\gamma$-ray source was detected with a significance of ${\sim}16\nobreak\,\sigma$ by KM2A and ${\sim}17\nobreak\,\sigma$ by WCDA, respectively. The angular extension of this $\gamma$-ray source is about 0.5 degrees, corresponding to a physical size of about 50 pc. We discuss the origin of the $\gamma$-ray emission and possible cosmic ray acceleration in the W43 region using multi-wavelength data. Our findings suggest that W43 is likely another young star cluster capable of accelerating cosmic rays (CRs) to at least several hundred TeV.
@@ 2.1 Data and analysis technique
Figure 1: These figures display significance maps of W43 derived from the best-fit elliptical model. All sources, except LHAASO J1848-0153u, are subtracted with their best-fit parameters. Figure (a) shows the significance map in the energy range from 25 TeV to 100 TeV. The green and red ‘+’ symbols indicate the best-fit positions of KM2A and WCDA, respectively, along with their corresponding $1\,\sigma$ statistical error bars. The solid ellipse represents the best-fit morphology of LHAASO J1848-0153u as determined by KM2A, while the dot-dash ellipse corresponds to that of WCDA. The parameters used to draw these elements are taken from the best-fit elliptical models with spectral parameters of GDE left free. Similarly, the blue ‘+’ symbol and the dashed circle show results from the Fermi data (Yang & Wang, 2020). The contour lines in black, dark green and bright green indicate gas column densities of $3$, $6$ and $10\times 10^{22}{\rm\ cm^{-2}}$ (Yang & Wang, 2020), respectively. Supernova Remnants (SNRs, marked by square symbols) from Green (2019) and pulsars (PSRs, marked by cross symbols) from ATNF pulsar catalogue (Manchester et al., 2005, 2023) are also shown here. The positions of the two molecular clouds W43-main and W43-south, the TeV source HESS J1848-018 and the globular cluster GLIMPSE C01 are also plotted. Figure (b) shows the significance map in the energy range from 1 TeV to 25 TeV. Figure (c) shows the significance map in the energy range above 100 TeV.
@@ 2.2 Results
After applying the aforementioned procedure, we derived the residual significance map. This map shows the residual significance after subtracting all sources except for LHAASO J1848-0153u. The subtraction process was performed in accordance with the best-fit model. This process unveiled excess emissions in the vicinity of W43, as depicted in Figure 1. The significance of signal excess reaches up to ${\sim}17\nobreak\,\sigma$ in the energy range from $1$ TeV to $25$ TeV, ${\sim}16\nobreak\,\sigma$ from $25$ TeV to $100$ TeV, and $3.6\ \sigma$ above $100~{}\rm TeV$.
@@ 3 DISCUSSIONS
Figure 2 shows the fit of the observed $\gamma$-ray spectra with the model (Kafexhiu et al., 2014) in which CRs interact with ambient gas, using NAIMA (Zabalza, 2015) package. In the fittings, we used the spectra derived in the previous section. We discovered that both a broken power law (BPL) and a power law with an exponential cutoff (ECPL) are suitable models for parent protons. The formulae for BPL and ECPL are shown in Table 3. The total CR energy budget ($>1\ {\rm TeV}$) is about $2.5\times 10^{48}~{}\rm erg$ assuming a gas density of about $140~{}\rm cm^{-3}$. The fitting results are shown in Table 3. We note that in this case a quite sharp cutoff of break/cutoff in proton spectra is needed to fit the $\gamma$-ray flux. The derived CR density, as shown in Figure 2 (b), is nearly 10 times higher than the CR density measured locally above 1 TeV.
@@
[…]
ECPL(PP) $2.27^{+0.05}_{-0.06}$ $3.10^{+1.80}_{-1.56}$ $406^{+142}_{-108}$ $2.51^{+0.42}_{-0.37}\times 10^{48}$ erg -4.58
BPL(PP) $2.27^{+0.05}_{-0.06}$ $4.41^{+0.92}_{-0.74}$ $211^{+96}_{-82}$ $2.51^{+0.40}_{-0.39}\times 10^{48}$ erg -4.00
ECPL(IC) $2.98^{+0.08}_{-0.09}$ $2.08^{+2.16}_{-1.21}$ $228^{+389}_{-87}$ $2.70^{+0.34}_{-0.32}\times 10^{47}$ erg -6.73

### v2

@@ Abstract
In this paper, we report the detection of the very-high-energy (VHE, $100{\rm\ GeV}<E<100{\rm\ TeV}$) and ultra-high-energy (UHE, $E>100\rm\ TeV$) $\gamma$-ray emissions from the direction of the young star-forming region W43, observed by the Large High Altitude Air Shower Observation (LHAASO). The extended $\gamma$-ray source was detected with a significance of ${\sim}16\penalty 10000\,\sigma$ by KM2A and ${\sim}17\penalty 10000\,\sigma$ by WCDA, respectively. The angular extension of this $\gamma$-ray source is about 0.5 degrees, corresponding to a physical size of about 50 pc. We discuss the origin of the $\gamma$-ray emission and possible cosmic ray acceleration in the W43 region using multi-wavelength data. Our findings suggest that W43 is likely another young star cluster capable of accelerating cosmic rays (CRs) to at least several hundred TeV.
@@ 2.1 Data and analysis technique
Figure 1: These figures display significance maps of W43 derived from the best-fit elliptical model. All sources, except LHAASO J1848-0153u, are subtracted with their best-fit parameters. Figure (a) shows the significance map in the energy range from 25 TeV to 100 TeV. The green and red ‘+’ symbols indicate the best-fit positions of KM2A and WCDA, respectively, along with their corresponding $1\,\sigma$ statistical error bars. The solid ellipse represents the best-fit morphology of LHAASO J1848-0153u as determined by KM2A, while the dot-dash ellipse corresponds to that of WCDA. The parameters used to draw these elements are taken from the best-fit elliptical models with spectral parameters of GDE left free. Similarly, the blue ‘+’ symbol and the dashed circle show results from the Fermi data (Yang & Wang, 2020). The contour lines in black, dark green and bright green indicate gas column densities of $3$, 66 and $10\times 10^{22}{\rm\ cm^{-2}}$ (Yang & Wang, 2020), respectively. Supernova Remnants (SNRs, marked by square symbols) from Green (2019) and pulsars (PSRs, marked by cross symbols) from ATNF pulsar catalogue (Manchester et al., 2005, 2023) are also shown here. The positions of the two molecular clouds W43-main and W43-south, the TeV source HESS J1848-018 and the globular cluster GLIMPSE C01 are also plotted. Figure (b) shows the significance map in the energy range from 1 TeV to 25 TeV. Figure (c) shows the significance map in the energy range above 100 TeV.
@@ 2.2 Results
After applying the aforementioned procedure, we derived the residual significance map. This map shows the residual significance after subtracting all sources except for LHAASO J1848-0153u. The subtraction process was performed in accordance with the best-fit model. This process unveiled excess emissions in the vicinity of W43, as depicted in Figure 1. The significance of signal excess reaches up to ${\sim}17\penalty 10000\,\sigma$ in the energy range from $1$ TeV to $25$ TeV, ${\sim}16\penalty 10000\,\sigma$ from $25$ TeV to $100$ TeV, and $3.6\ \sigma$ above $100~\rm TeV$.
@@ 3 DISCUSSIONS
Figure 2 shows the fit of the observed $\gamma$-ray spectra with the model (Kafexhiu et al., 2014) in which CRs interact with ambient gas, using NAIMA (Zabalza, 2015) package. In the fittings, we used the spectra derived in the previous section. We discovered that both a broken power law (BPL) and a power law with an exponential cutoff (ECPL) are suitable models for parent protons. The formulae for BPL and ECPL are shown in Table 3. The total CR energy budget ($1\ {\rm TeV}$) is about $2.5\times 10^{48}~\rm erg$ assuming a gas density of about $140~\rm cm^{-3}$. The fitting results are shown in Table 3. We note that in this case a quite sharp cutoff of break/cutoff in proton spectra is needed to fit the $\gamma$-ray flux. The derived CR density, as shown in Figure 2 (b), is nearly 10 times higher than the CR density measured locally above 1 TeV.
@@
Distribution $\alpha/\alpha_{1}$ $\beta/\alpha_{2}$ $E_{\rm cut}/E_{\rm b}$ $W_{p}(>1~\rm TeV){\textsuperscript{2)}}$ ${\rm MLL}{\textsuperscript{3)}}$
ECPL($P$) $2.27^{+0.05}_{-0.06}$ $3.10^{+1.80}_{-1.56}$ $406^{+142}_{-108}$ $2.51^{+0.42}_{-0.37}\times 10^{48}$ erg -4.58
BPL($P$) $2.27^{+0.05}_{-0.06}$ $4.41^{+0.92}_{-0.74}$ $211^{+96}_{-82}$ $2.51^{+0.40}_{-0.39}\times 10^{48}$ erg -4.00
ECPL(IC) $2.98^{+0.08}_{-0.09}$ $08^{+2.16}_{-1.21} 22$$28^{+389}_{-87}$ $2.70^{+0.34}_{-0.32}\times 10^{47}$ erg -6.73